# SeNMo: A Self-Normalizing Deep Learning Model for Enhanced Multi-Omics Data Analysis in Oncology

### v1

@@ 2.1.4 Features integration
After carrying out the preprocessing steps mentioned above, we integrate the data across cancers and across modalities. We generate two views of the data by concatenating the features across cancers and across modalities. First view is created by taking the union of features across all cancer patients for each of the six modalities (DNA methylation, gene expression, miRNA expression, protein expression, DNA mutation, and clinical). The DNA methylation data originally had 485,576 features across all cancers, which was reduced to $\approx$ 4500 features. Union of these features resulted in the feature dimension of 52,396. The gene expression data originally had 60,483 features across all cancers, which was reduced to $\approx$ 300 features. Union of these features resulted in the feature dimension of 8794. The miRNA expression data originally had 1880 features across all cancers, which was reduced to $\approx$ 1400 features. Union of these features resulted in the feature dimension of 1730. The protein expression data originally had 487 features across all cancers, which was reduced to 472 features unionized to 472 dimensions. The DNA mutation data had 18,090 features across all cancers, pre-processed and unionized to 17,253 features. Lastly, we convert the categorical clinical features to numerical values such as gender, race, and cancer stages. The details of these clinical characteristics are given in Table 2. Mathematically, the preprocessing is given below.
Let $D$ represent the initial feature having fixed dimension for each cancer. The dimension of each feature set is reduced through a preprocessing step to $\tilde{D}$, which is calculated by a function of $D$, noted as $f(D)$, where $f$ is the dimension reduction function such as those presented in the previous section, $\tilde{D}=f(D)$. The reduced dimensions $\tilde{D}$ from each cancer type are then combined through a union operation to generate a feature vector $D_{m}$ for each modality $m$ and $M=6$ are the total number of modalities. For $n=33$ cancer types and $\tilde{D}$ is the preprocessed, reduced dimension feature vector, the $D_{m}$ is the union of these dimensions. Finally, the union of all $D_{m}$ across different modalities results in the total pan-cancer, multimodal feature vector dimension $D_{c}\in\mathbb{R}^{80,697}$. The feature vector dimension for each modality, $D_{m}$, is defined as:
$D_{m}=\begin{cases}\bigcup_{i=1}^{n}\tilde{D}_{i}&\text{if }\tilde{D}_{i}\text{ varies by cancer type or modality,}\\ \tilde{D}&\text{otherwise.}\end{cases}$ (1)
The total pan-cancer, multimodal feature vector dimension $D_{c}$ can then be expressed as:
$D_{c}=\bigcup_{m=1}^{M}D_{m}$ (2)
@@
Cancer Type No. of Cases Age (Mean±SD) Gender (M/F) Race (White/Asian/Black/NA/American Indian/Alaska/Islander) Stage (0/I/IA/IB/IC/II/IIA/IIB/IIC/III/IIIA/IIIB/IIIC/IV/IVA/IVB/IVC/NA)
TCGA-ACC 95 47.46 ± 16.2 33/62 79/3/1/12/0/0 0/9/0/0/0/46/0/0/0/20/0/0/0/17/0/0/0/3
TCGA-BLCA 447 67.92 ± 10.39 326/121 363/43/23/18/0/0 0/3/0/0/0/136/0/0/0/159/0/0/0/148/0/0/0/1
TCGA-BRCA 1260 57.94 ± 13.11 13/1247 915/59/198/87/1/0 0/114/94/7/0/6/404/307/0/2/176/30/74/22/0/0/0/24
TCGA-CESC 304 48.04 ± 13.7 0/304 211/19/32/30/9/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/304
TCGA-CHOL 62 64.37 ± 12.21 30/32 55/3/3/1/0/0 0/30/0/0/0/16/0/0/0/5/0/0/0/2/3/6/0/0
TCGA-COAD 539 66.93 ± 12.67 288/251 261/11/67/198/2/0 0/87/1/0/0/46/150/13/2/26/9/69/47/56/18/3/0/12
TCGA-DLBC 51 56.76 ± 13.68 24/27 32/18/1/0/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/51
TCGA-ESCA 249 64.22 ± 12.11 208/41 162/46/6/35/0/0 0/14/9/7/0/1/56/43/0/41/16/10/9/7/6/0/0/30
TCGA-GBM 649 57.74 ± 14.32 399/250 547/13/53/36/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/649
TCGA-HNSC 611 61.02 ± 11.92 443/168 522/12/58/17/2/0 0/29/0/0/0/93/0/0/0/97/0/0/0/0/302/13/1/76
TCGA-KICH 182 51.61 ± 14.12 99/83 154/6/19/3/0/0 0/75/0/0/0/59/0/0/0/34/0/0/0/14/0/0/0/0
TCGA-KIRC 979 60.67 ± 11.95 641/338 876/16/73/14/0/0 0/475/0/0/0/102/0/0/0/237/0/0/0/161/0/0/0/4
TCGA-KIRP 376 61.98 ± 12.2 278/98 275/6/75/16/4/0 0/219/0/0/0/25/0/0/0/77/0/0/0/21/0/0/0/34
TCGA-LAML 626 54.82 ± 15.87 345/281 564/8/49/5/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/626
TCGA-LGG 533 42.71 ± 13.32 293/240 492/8/22/10/1/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/533
TCGA-LIHC 463 60.44 ± 13.71 305/158 255/168/25/14/1/0 0/211/0/0/0/105/0/0/0/6/78/12/11/2/1/3/0/34
TCGA-LUAD 728 65.20 ± 10.08 329/399 580/14/84/48/2/0 0/7/194/195/0/2/67/103/0/0/101/12/0/37/0/0/0/10
TCGA-LUSC 752 67.28 ± 8.62 548/204 530/12/47/163/0/0 0/4/127/243/0/4/87/138/0/3/94/33/0/12/0/0/0/7
TCGA-MESO 86 63.01 ± 9.78 70/16 84/1/1/0/0/0 0/7/2/1/0/15/0/0/0/45/0/0/0/16/0/0/0/0
TCGA-OV 731 59.60 ± 11.44 0/731 626/25/43/33/3/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/731
TCGA-PAAD 222 64.87 ± 11.36 123/99 195/13/8/6/0/0 0/1/6/15/0/0/36/148/0/6/0/0/0/7/0/0/0/3
TCGA-PCPG 189 47.02 ± 15.15 84/105 157/7/20/4/1/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/189
TCGA-PRAD 623 60.93 ± 6.8 623/0 510/13/81/18/1/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/623
TCGA-READ 178 63.83 ± 11.85 98/80 90/1/7/80/0/0 0/37/0/0/0/7/40/2/1/6/7/25/14/21/7/0/0/11
TCGA-SARC 287 60.70 ± 14.38 129/158 253/5/20/9/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/287
TCGA-SKCM 436 57.84 ± 15.41 289/174 441/12/1/9/0/0 6/30/18/30/0/39/18/28/61/44/16/46/68/23/0/0/0/36
TCGA-STAD 499 65.44 ± 10.53 320/179 311/108/15/64/0/0 0/1/21/46/0/37/54/71/0/4/88/67/39/47/0/0/0/24
TCGA-TGCT 139 31.87 ± 9.19 139/0 124/4/6/5/0/0 0/69/26/11/0/4/6/1/1/2/1/6/5/0/0/0/0/7
TCGA-THCA 614 47.17 ± 15.83 166/448 413/59/35/106/1/0 0/350/0/0/0/64/0/0/0/134/0/0/0/4/52/0/8/2
TCGA-THYM 138 58.12 ± 13 72/66 115/13/8/2/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/138
TCGA-UCEC 588 63.74 ± 11.06 0/588 402/21/120/32/4/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/588
TCGA-UCS 61 70.07 ± 9.24 0/61 50/1/9/1/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/61
TCGA-UVM 80 61.65 ± 13.95 45/35 55/0/0/25/0/0 0/0/0/0/0/0/12/27/0/0/25/10/1/4/0/0/0/1
Moffitt-LSCC 108 69.14 ± 8.34 72/36 105/0/3/0/0/0 0/0/24/25/0/0/31/15/0/0/12/1/0/0/0/0/0/0
Figure 2: Features processing pipeline for pan-cancer data.
@@ 2.2 Clinical end-points
To assess the performance of the SeNMo framework, we chose two end-points that belong to two different categories of machine learning tasks. First is the clinically relevant prediction of overall survival (OS) which is a regression task. The second is the prediction of the primary cancer type which is 33-class classification task.
@@ 2.2.1 Overall Survival (OS)
Cancer prognosis via survival outcome prediction is a standard method used for biomarker discovery, stratification of patients into distinct treatment groups, and therapeutic response prediction [102]. Advancements in next-generation sequencing have led to a surge in data availability [28]. Statistical survival models and a shift towards incorporating deep learning into survival analysis has improved the task of OS prediction. Previous works have used combinations of molecular data types and employed different statistical and learning-based methods to predict OS in different datasets [103, 33, 35, 36]. The journey towards comprehensive survival analysis continues, aiming to combine different data types to better understand the relationship between molecular features and patient outcomes, ultimately leading to more precise prognostic assessments and tailored treatment approaches. In this study, we utilize clinical, demographic, genomic, and molecular data types to investigate potential risk factors for cancer patients, specifically examining their correlations with the patients’ time-to-event, which in this case is overall survival. We have implemented prediction of OS as a regression task, i.e., prediction of OS in days. Time-to-event or survival data not only records the occurrence of events like death but also tracks the time from the study’s outset to when the event happens, tracks when the study concludes, or tracks when a patient is no longer followed (known as right censoring). Survival or right-censored times since cancer diagnosis for our pan-cancer data is depicted in Figure 3A. Due to censoring, the exact survival time for some patients remains unknown. In such instances, each patient’s outcome is defined by two main variables: a censoring indicator, also called the status, and the observed time $T=\min(T_{s},T_{c})$, where ${T}_{s}$ represents the exact survival time and ${T}_{c}$ is the censoring time, $\{{T}_{s}\leq{T}_{c}\}$ [10]. Survival function to describe the likelihood of a patient surviving beyond a specified time $t$, is given as:
@@
Additionally, the hazard function offers insights into the risk of the event occurring at a given time, assuming survival up to that point. Hazard function represents the instantaneous rate at which events (such as death) occur at a specific time, given that the individual has survived up to that time. It provides insights into the risk of experiencing the event at any given moment, conditional on having survived up to that point in time. Mathematically, the hazard function $h(t)$ is defined as the ratio of the probability of the event occurring in a small time interval around $t$ to the probability of surviving beyond $t$:
Figure 3: Summary of number of cases in the pan-cancer data.
$h(t)=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t\,|\,T\geq t)}{\Delta t}$ (4)
where:
  • $h(t)$ is the hazard function at time $t$.
  • $T$ is the survival time.
  • $P(t\leq T<t+\Delta{t}\,|\,T\geq{t})$ is the conditional probability that the event occurs in the time interval $[t,t+\Delta{t})$ given that survival time is greater than or equal to $t$.
  • $\Delta{t}$ represents an infinitesimally small time interval.
Hazard function describes the instantaneous risk of experiencing the event of interest at any given time, based on survival data. In our pan-cancer data, the (right)censoring was defined as $censor=1$ in case of an event (death), and 00 otherwise.
@@ 2.2.2 Primary Cancer type
The primary cancer type prediction task involves classifying a cancer sample into one of 33 possible types based on various biological and clinical features. This task is fundamentally important and clinically relevant because accurate identification of the primary cancer type is critical for choosing the most effective treatment strategy, improving patient outcomes, and personalizing therapy approaches [104]. Cancer treatments and prognostics can vary dramatically between different cancer types, often requiring specific interventions that are tailored to the unique biological characteristics of each type. Correctly predicting the primary cancer type helps in planning follow-up care and surveillance, enhancing the likelihood of early detection of recurrence. Thus, achieving high accuracy in this classification not only supports better clinical decision-making but also significantly impacts patient survival and quality of life. The pan-cancer data comprising of 33 cancers and distribution of patients (cases) across these datasets are depicted in Figure 3B.

### v2

@@ 2.1.4 Features integration
After carrying out the preprocessing steps mentioned above, we integrate the data across cancers and across modalities. We generate two views of the data by combining the features across cancers and across modalities. First view is created by taking the union of features across all cancer patients for each of the six modalities (DNA methylation, gene expression, miRNA expression, protein expression, DNA mutation, and clinical). As a result of the preprocessing explained earlier, the DNA methylation data features were reduced from 485,576 features to $\approx$ 4,500 features for all cancers. The union of these features from individual cancers resulted in a feature dimension of 52,396. The gene expression data originally had 60,483 features across all cancers, which was reduced to $\approx$ 3000 features. Union of these features resulted in the feature dimension of 8,794. The miRNA expression data originally had 1,880 features across all cancers, which was reduced to $\approx$ 1,400 features. Union of these features resulted in the feature dimension of 1,730. The protein expression data originally had 487 features across all cancers, which was reduced to 472 features unionized to 472 dimensions. The DNA mutation data had 18,090 features across all cancers, pre-processed and unionized to 17,253 features. Lastly, we convert the categorical clinical features to numerical values such as gender, race, and cancer stages. The details of these clinical characteristics are given in Table 2. Mathematically, the preprocessing is given below.
Let v represent the initial feature having fixed dimension for each cancer. The dimension of each feature set is reduced through a preprocessing step, resulting in the feature vector $\tilde{\textbf{v}}$, which is calculated by a function of v, noted as $f(\textbf{v})$, where $f$ is the dimension reduction function such as those presented in the previous section, $=f(\textbf{v})$ v ). For $n=33$ cancer types, the reduced dimensional feature vector $\tilde{\textbf{v}}$ from each cancer type are then combined through a union operation to generate a feature vector $V_{m}$ for each modality $m$ and $M=6$ are the total number of modalities. The feature vector for each modality, $V_{m}$, is defined as:
$V_{m}=\begin{cases}\bigcup_{i=1}^{n}\tilde{\textbf{v}}_{i}&\text{if }\tilde{\textbf{v}}_{i}\text{ varies by cancer type or modality,}\\ \tilde{\textbf{v}}&\text{otherwise.}\end{cases}$ (1)
Finally, the union of all $V_{m}$ across different modalities results in the total pan-cancer, multimodal feature vector $V_{c}\in\mathbb{R}^{80,697}$. The total pan-cancer, multimodal feature vector $V_{c}$ can then be expressed as:
$V_{c}=\bigcup_{m=1}^{M}V_{m}$ (2)
@@
Cancer Type Age (Mean±SD) Gender (M/F) Race (White/Asian/Black/NA/American Indian/Alaska/Islander) Stage (0/I/IA/IB/IC/II/IIA/IIB/IIC/III/IIIA/IIIB/IIIC/IV/IVA/IVB/IVC/NA)
TCGA-ACC 47.46 ± 16.2 33/62 79/3/1/12/0/0 0/9/0/0/0/46/0/0/0/20/0/0/0/17/0/0/0/3
TCGA-BLCA 67.92 ± 10.39 326/121 363/43/23/18/0/0 0/3/0/0/0/136/0/0/0/159/0/0/0/148/0/0/0/1
TCGA-BRCA 57.94 ± 13.11 13/1247 915/59/198/87/1/0 0/114/94/7/0/6/404/307/0/2/176/30/74/22/0/0/0/24
TCGA-CESC 48.04 ± 13.7 0/304 211/19/32/30/9/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/304
TCGA-CHOL 64.37 ± 12.21 30/32 55/3/3/1/0/0 0/30/0/0/0/16/0/0/0/5/0/0/0/2/3/6/0/0
TCGA-COAD 66.93 ± 12.67 288/251 261/11/67/198/2/0 0/87/1/0/0/46/150/13/2/26/9/69/47/56/18/3/0/12
TCGA-DLBC 56.76 ± 13.68 24/27 32/18/1/0/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/51
TCGA-ESCA 64.22 ± 12.11 208/41 162/46/6/35/0/0 0/14/9/7/0/1/56/43/0/41/16/10/9/7/6/0/0/30
TCGA-GBM 57.74 ± 14.32 399/250 547/13/53/36/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/649
TCGA-HNSC 61.02 ± 11.92 443/168 522/12/58/17/2/0 0/29/0/0/0/93/0/0/0/97/0/0/0/0/302/13/1/76
TCGA-KICH 51.61 ± 14.12 99/83 154/6/19/3/0/0 0/75/0/0/0/59/0/0/0/34/0/0/0/14/0/0/0/0
TCGA-KIRC 60.67 ± 11.95 641/338 876/16/73/14/0/0 0/475/0/0/0/102/0/0/0/237/0/0/0/161/0/0/0/4
TCGA-KIRP 61.98 ± 12.2 278/98 275/6/75/16/4/0 0/219/0/0/0/25/0/0/0/77/0/0/0/21/0/0/0/34
TCGA-LAML 54.82 ± 15.87 345/281 564/8/49/5/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/626
TCGA-LGG 42.71 ± 13.32 293/240 492/8/22/10/1/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/533
TCGA-LIHC 60.44 ± 13.71 305/158 255/168/25/14/1/0 0/211/0/0/0/105/0/0/0/6/78/12/11/2/1/3/0/34
TCGA-LUAD 65.20 ± 10.08 329/399 580/14/84/48/2/0 0/7/194/195/0/2/67/103/0/0/101/12/0/37/0/0/0/10
TCGA-LUSC 67.28 ± 8.62 548/204 530/12/47/163/0/0 0/4/127/243/0/4/87/138/0/3/94/33/0/12/0/0/0/7
TCGA-MESO 63.01 ± 9.78 70/16 84/1/1/0/0/0 0/7/2/1/0/15/0/0/0/45/0/0/0/16/0/0/0/0
TCGA-OV 59.60 ± 11.44 0/731 626/25/43/33/3/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/731
TCGA-PAAD 64.87 ± 11.36 123/99 195/13/8/6/0/0 0/1/6/15/0/0/36/148/0/6/0/0/0/7/0/0/0/3
TCGA-PCPG 47.02 ± 15.15 84/105 157/7/20/4/1/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/189
TCGA-PRAD 60.93 ± 6.8 623/0 510/13/81/18/1/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/623
TCGA-READ 63.83 ± 11.85 98/80 90/1/7/80/0/0 0/37/0/0/0/7/40/2/1/6/7/25/14/21/7/0/0/11
TCGA-SARC 60.70 ± 14.38 129/158 253/5/20/9/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/287
TCGA-SKCM 57.84 ± 15.41 289/174 441/12/1/9/0/0 6/30/18/30/0/39/18/28/61/44/16/46/68/23/0/0/0/36
TCGA-STAD 65.44 ± 10.53 320/179 311/108/15/64/0/0 0/1/21/46/0/37/54/71/0/4/88/67/39/47/0/0/0/24
TCGA-TGCT 31.87 ± 9.19 139/0 124/4/6/5/0/0 0/69/26/11/0/4/6/1/1/2/1/6/5/0/0/0/0/7
TCGA-THCA 47.17 ± 15.83 166/448 413/59/35/106/1/0 0/350/0/0/0/64/0/0/0/134/0/0/0/4/52/0/8/2
TCGA-THYM 58.12 ± 13 72/66 115/13/8/2/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/138
TCGA-UCEC 63.74 ± 11.06 0/588 402/21/120/32/4/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/588
TCGA-UCS 70.07 ± 9.24 0/61 50/1/9/1/0/0 0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/0/61
TCGA-UVM 61.65 ± 13.95 45/35 55/0/0/25/0/0 0/0/0/0/0/0/12/27/0/0/25/10/1/4/0/0/0/1
Moffitt-LSCC 69.14 ± 8.34 72/36 105/0/3/0/0/0 0/0/24/25/0/0/31/15/0/0/12/1/0/0/0/0/0/0
Figure 2: Features processing pipeline for pan-cancer data encompassing six data types: DNA Methylation, Gene Expression, miRNA Expression, Protein Expression, DNA Mutation, and Clinical features. Initial feature counts are reduced through preprocessing, with each modality unified at the pan-cancer level to yield unimodal pan-cancer feature sets. These unimodal features are further unified across all modalities, resulting in a final integrated multimodal feature matrix of 80,697 features, used for downstream analysis.
@@ 2.2 Clinical end-points
To assess the performance of the SeNMo framework, we selected three clinical end-points that fall under two categories of machine learning tasks. The first end-point is Overall Survival (OS), which is treated as a regression task. The second is the prediction of primary cancer type, formulated as a 33-class classification task. The third end-point is TLS ratio prediction, also a regression task.
@@ 2.2.1 Overall Survival (OS)
Predicting cancer prognosis through survival outcomes is a standard approach for biomarker discovery, patient stratification, and assessing therapeutic response [101]. Statistical survival models, coupled with the integration of deep learning in survival analysis, have significantly advanced the prediction of OS. Prior studies have combined different molecular data types and employed a range of statistical and machine learning methods to predict OS across various datasets [102, 34, 36, 37]. This ongoing effort aims to integrate multiple data types to elucidate the relationship between molecular characteristics and patient outcomes, ultimately achieving more precise prognostic assessments and personalized treatment strategies. In this study, we utilize clinical, demographic, genomic, and other molecular data to explore potential risk factors for cancer patients and to analyze their correlation with the patients’ time-to-event, specifically OS. The prediction of OS is implemented as a regression task, with the goal of predicting survival time in days. Time-to-event or survival data records not only the occurrence of events such as death but also the duration from the beginning of the study until the event occurs, or until the patient is lost to follow-up (right censoring). Survival times since cancer diagnosis for the pan-cancer dataset are depicted in Figure 3A. Because of censoring, exact survival times are unknown for some patients. In these cases, each patient’s outcome is characterized by two variables: a censoring indicator, also known as the vital status, and the observed time $T=\min(T_{s},T_{\delta})$, where ${T}_{s}$ represents the true survival time and ${T}_{\delta}$ is the censoring time, $\{{T}_{s}\leq{T}_{\delta}\}$ [10]. The survival function, which describes the probability that a patient will survive beyond a specified time $t$, is given by:
@@
Additionally, the hazard function provides insight into the risk of an event occurring at a particular time, given survival up to that point. It represents the instantaneous rate of events (e.g., death) occurring at a specific time, conditional on having survived to that time. The hazard function $h(t)$ is mathematically defined as the ratio of the probability of the event occurring in a short interval around $t$ to the probability of surviving beyond $t$:
$h(t)=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t\,|\,T\geq t)}{\Delta t},$ (4)
where, $h(t)$ is the hazard function at time $t$, $T$ is the survival time, $P(t\leq T<t+\Delta{t}\,|\,T\geq{t})$ is the conditional probability that the event occurs in the time interval $[t,t+\Delta{t})$ given that survival time is greater than or equal to $t$, and $\Delta{t}$ represents an infinitesimally small time interval. Based on survival data, the hazard function describes the instantaneous risk of experiencing the event of interest at any given time. In our study, right-censoring was defined as censor $\delta=1$ in case of an event (e.g., death), and 00 otherwise.
Figure 3: Summary of the survival data and number of cases in the pan-cancer data. The top panel illustrates patient survival in years, with each line representing an individual patient, marked as censored or dead, and grouped by cancer type. The bottom panel displays the number of cases for each primary cancer type, highlighting variation in sample sizes across cancers, which range from high counts in breast, kidney, and lung to lower counts in rare cancers like diffuse large B-cell lymphoma.
@@ 2.2.2 Primary Cancer type
The prediction of the primary cancer type involves classifying each cancer sample into one of 33 possible cancer types based on biological and clinical features. This classification task is crucial for clinical decision-making, as accurate identification of the primary cancer type is essential for determining the most effective treatment approach, thereby improving patient outcomes and enabling personalized therapies [103]. Cancer treatments and prognoses differ significantly across cancer types, often necessitating specific, tailored interventions that align with the distinct biological characteristics of each type. Correct identification of the primary cancer type also aids in follow-up care and surveillance, increasing the likelihood of early detection of recurrence. Therefore, achieving high accuracy in this classification not only enhances clinical decision-making but also positively impacts patient survival and quality of life. The pan-cancer dataset encompassing 33 cancer types, along with the distribution of patient samples, is depicted in Figure 3.
2.2.3 Tertiary Lymphoid Structures (TLS) Ratio
TLSs are organized accumulations of immune cells that resemble secondary lymphoid organs and form in inflamed peripheral tissues, including within cancers [104, 105]. TLS presence is linked to improved survival rates and favorable responses to immunotherapy across various solid tumors, making TLS quantification a promising predictive and prognostic biomarker [104, 105]. The TLS ratio, defined as the segmented TLS area over the total tissue area, is correlated with positive immunotherapy outcomes and overall patient prognosis. Recent studies have demonstrated the value of TLS quantification in cancer, highlighting its role in improving clinical decision-making and developing automated TLS segmentation models with high accuracy in multiple cancers [104, 105]. In this study, we used whole slide images of H&E and CD20-stained sections imported into Visiopharm software version 2022.03. Visiopharm’s Tissuealign tool was used to co-register serial H&E and CD20 images for each patient. Using the H&E image, manually drawn regions of interest (ROIs) were created to segment the tumor and non-tumor regions in each image set. TLSs were detected through a thresholding algorithm, followed by manual review and feature extraction for analysis by an experienced image analysis technician under the guidance of the study pathologist.
Figure 4: Architecture of the SeNMo encoder network. There are seven hidden layers each comprising of a linear unit, SELU activation, and alpha-dropout. The trained model has 83.33 million parameters. The number of neurons in each hidden layer, input layer, and output layer are also depicted in the figure. The same model is used for regression and classification tasks.